# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

### 12 Jan 2024 v1

@@ IV-C3 Metrics
  • Success Rate (SuccRate): As repository-level dependencies can potentially introduce dependency errors in generated functions, we introduce the Success Rate metric (SuccRate) to evaluate the effectiveness of ToolGen in reducing dependency errors. This metric evaluates the proportion of generated functions that successfully pass a static check for dependency errors, specifically no-member and undefined-variable. To perform this evaluation, we incorporate the generated functions into their respective code repositories and conduct static lint analysis using pylint [30]. Functions that do not exhibit syntax errors, no-member, or undefined-variable errors are deemed successful. It’s worth noting that this work does not address logic errors, as their identification typically necessitates the execution of corresponding unit tests.
@@ V-A RQ1: Similarity Metrics
TABLE III: Evaluation results for repository-level metrics. DepCov and SuccRate represent Dependency Coverage and Success Rate, respectively. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
@@
V-C RQ2.b: Success Rate
Table III also presents the evaluation results for success rate (SuccRate) in repository-level lint analysis, with a particular focus on no-member and undefined-variable errors. Remarkably, our approach, ToolGen, consistently exhibits significantly higher success rates compared to the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen increases the success rate by 10.9%, 42.2%, and 40.4%, respectively.
@@
SUMMARY: Our approach, ToolGen, consistently achieves significantly higher success rates in repository-level lint analysis compared to the baselines, with improvements ranging from 10.9% to 42.2%. These results underscore the effectiveness of our tool-integrated generation process in mitigating the generation of invalid identifiers, a common challenge faced by conventional code LLMs in the context of repository-level code generation.
@@ V-D RQ3: Generalizability
Specifically, for the encoder-decoder model CodeT5, our tool-integrated generation approach demonstrates well-balanced and comprehensive enhancements in repository-level metrics. However, for the two decoder-only models, CodeGPT and CodeLlama, the improvements achieved by ToolGen vary between dependency coverage and success rate. When employing the smaller CodeGPT as the base model, ToolGen significantly enhances dependency coverage but exhibits less pronounced improvements in the success rate. In contrast, when employing the larger CodeLlama, the results are reversed. Upon closer examination of the generated functions, we attribute these variations to CodeGPT’s occasional difficulty in predicting the <EOS> token correctly, leading to syntax errors that violate the prerequisites for successful lint check. In the case of CodeLlama, the initially high dependency coverage, owing to its extensive parameter scale and training data, poses challenges for improvement through parameter-efficient fine-tuning (PEFT) with LoRA (only 0.25% trainable parameters).

### 21 Jan 2024 v2

@@ IV-C3 Metrics
  • Static Validity Rate (ValRate): As repository-level dependencies can potentially introduce dependency errors in generated functions, we introduce the Static Validity Rate metric (ValRate) to evaluate the effectiveness of ToolGen in reducing dependency errors. This metric evaluates the proportion of generated functions that successfully pass a static check for dependency errors, specifically no-member and undefined-variable. To perform this evaluation, we incorporate the generated functions into their respective code repositories and conduct static lint analysis using pylint [30]. Functions that do not exhibit syntax errors, no-member, or undefined-variable errors are deemed statically valid. It’s worth noting that this work does not address logic errors, as their identification typically necessitates the execution of corresponding unit tests.
@@ V-A RQ1: Similarity Metrics
TABLE III: Evaluation results for repository-level metrics. DepCov and ValRate represent Dependency Coverage and Static Validity Rate, respectively. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
@@
V-C RQ2.b: Static Validity Rate
Table III also presents the evaluation results for Static Validity Rate (ValRate) in repository-level lint analysis, with a particular focus on no-member and undefined-variable errors. Remarkably, our approach, ToolGen, consistently exhibits significantly higher Static Validity Rate compared to the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen increases the Static Validity Rate by 10.9%, 42.2%, and 40.4%, respectively.
@@
SUMMARY: Our approach, ToolGen, consistently achieves significantly higher Static Validity Rate in repository-level lint analysis compared to the baselines, with improvements ranging from 10.9% to 42.2%. These results underscore the effectiveness of our tool-integrated generation process in mitigating the generation of invalid identifiers, a common challenge faced by conventional code LLMs in the context of repository-level code generation.
@@ V-D RQ3: Generalizability
Specifically, for the encoder-decoder model CodeT5, our tool-integrated generation approach demonstrates well-balanced and comprehensive enhancements in repository-level metrics. However, for the two decoder-only models, CodeGPT and CodeLlama, the improvements achieved by ToolGen vary between Dependency Coverage and Static Validity Rate. When employing the smaller CodeGPT as the base model, ToolGen significantly enhances Dependency Coverage but exhibits less pronounced improvements in the Static Validity Rate. In contrast, when employing the larger CodeLlama, the results are reversed. Upon closer examination of the generated functions, we attribute these variations to CodeGPT’s occasional difficulty in predicting the <EOS> token correctly, leading to syntax errors that violate the prerequisites for successful lint check. In the case of CodeLlama, the initially high Dependency Coverage, owing to its extensive parameter scale and training data, poses challenges for improvement through parameter-efficient fine-tuning (PEFT) with LoRA (only 0.25% trainable parameters).